24. A cylinder has a stroke length of 1.6 meters. The time for th entire stroke is 12 seconds. What average velocity must the The time piston to move through the piston attain? 25. If the piston in the cylinder of problem 24 has a diameter of four rate (in GPM) is required to produce the desired velocity? moves a 7.5-ton load, what pressure will be required? 26. If the piston in problem 25 27, Thirty-eight liters per minute is flowing in a 3 ¾ cm ID pipe. What will the fluid velocity be in cm per second? 28. If the pipe in problem 27 is increased to a 7 ½ cm ID. pipe (diameter is doubled), will the velocity of the fluid increase or decrease? By how much? 29. Four horsepower is converted to heat for 12 seconds. How many BTUs does this represent?
Answer #1

24) Given, L=1.6 m, t= 12s, average velocity V = L/t=1.6/12=0.1333 m/s

25)Given d= 4 inches=0.1016 m, flow rate Q= Area×Velocity=pi xd2/4 x V = 0.0081 m3/s=128.387 GPM (1 m3/s =15,850.32 GPM)

26) Given F=7.5 ton = 73549.875 Newtons, F= pressure × Area=p ×0.0081

P=73549.875/0.0081=9080231.481N/m2=90.802 bar

26)